peedy Couriers documented the miles driven and total vehicle costs for the past five months as

follows:

Number of miles

Total vehicle costs

January

1,600

$3,600

February

2,000

4,600

March

1,500

3,450

April

1,800

3,900

May

2,200

4,500

Refer to the Speedy Couriers information above. Using the high/low method, if Speedy expects to drive 1,750 miles in June, what will be expected total vehicle costs?  (Select closest number.)

$3,825.00

$4,585.00

$8,087.50

$3,855.00
Answer #1

Variable cost per mile under High low method = (Total vehicle cost under highest point - Total vehilce cost under lowest point) / Difference in highest lowest point

Highest point = 2200 units $4500

Lowest point = 1500 units $3450

= (4500 - 3450) / (2200 - 1500)

= 1050 / 700

= $1.50 per mile

Fixed cost = 3450 - 1500 * 1.50 = 3450 - 2250

= 1200

Total cost with 1750 miles = Total miles * Variable cost per mile + Fixed cost

1750 * 1.50 + 1200

= 2625 + 1200

= 3825